OpenStudy (ksaimouli):

y=cos2u u=3x+1 find derivative

OpenStudy (ksaimouli):

chain rule

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}\]

OpenStudy (hba):

Its An Easy Question First Differentiate both with respect to u ^^

OpenStudy (anonymous):

\[\frac{dy}{du}=-\cos(u)\] and \[\frac{du}{dx}=3\]

OpenStudy (anonymous):

oops \[\frac{dy}{du}=-\sin(x)\]

OpenStudy (ksaimouli):

-6sinx

OpenStudy (ksaimouli):